OpenStudy (nuccioreggie):

How many solutions does the equation 6a − 3a − 6 = −2 + 3 have? Two None One Infinitely many

OpenStudy (nuccioreggie):

@whovianchick

OpenStudy (whovianchick):

6a - 3a - 6 = -2 + 3 3a - 6 = 1 3a = 7 There would only be one solution.